Circular No. 5916

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 A. D. Scott and A. Evans, University of Keele; S. Rawlings, Oxford

University; and S. Eales, University of Toronto, communicate: “We obtained a K-band spectrum of Nova Cas 1993 with the cooled grating spectrometer CGS4 on the U.K. Infrared Telescope on 1993 Dec. 13. In addition to Na I (2.21 μm), Brγ (2.16 μm), and C I (2.12 μm) in emission, and possibly He I (2.06 μm) in absorption, the spectrum also shows strong emission in the first overtone vibrational-rotational bands of ^{12}CO, up to at least the v = 6→4 transition; there may also be some evidence for ^{13}CO.”

PERIODIC COMET KUSHIDA-MURAMATSU (1993t) Improved orbital elements from MPEC 1994-A03:

T = 1993 Dec. 6.731 TT  ω = 347.^{o}499
e = 0.27592  Ω = 93.670 }2000.0
q = 2.75036 AU  i = 2.379

a = 3.79842 AU n^{o} = 0.133137 P = 7.40 years

1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 8 5^{h}13.^{m}62 +22^{o}43.^{′}4 1.850 2.756 151.^{o}8 9.^{o}7 16.7
18 5 09.60 +22 48.1 1.928 2.760 140.7 13.1 16.8
28 5 08.19 +22 55.0 2.025 2.765 130.2 15.8 16.9
Feb. 7 5 09.49 +23 04.4 2.139 2.771 120.3 17.9 17.1
17 5 13.39 +23 15.8 2.265 2.778 111.1 19.4 17.2
27 5 19.69 +23 28.3 2.399 2.786 102.5 20.3 17.4

NOTICE TO SUBSCRIBERS Subscribers to the Computer Service will notice some additional fea-

tures when next they log in. There is now the possibility of reading (using options ‘O’ and ‘P’) older IAU Circulars and Minor Planet Electronic Circulars long after they have been moved from the current files; the available IAUCs in fact go back to 1982, and there is a complete index all the way back to IAUC 1 in 1922. Option ‘A’ contains other new features, including information about specific supernovae and novae, and there is a current critical list for observations of and discovery information about numbered minor planets.

1994 January 6 Daniel W. E. Green